OpenStudy (anonymous):

In a multiple choice there are 20 questions. for each question there is a choice of four answers and only of of these is correct. Suppose 5 students guess answers to the test. What is the probability that at least two of them get more than ten answers correct?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how far did you get here?

jimthompson5910 (jim_thompson5910):

what are n and p in this case?

OpenStudy (anonymous):

probability of correct = 0.25... is this p? and q=20? i have no idea. I only understand that the question seems to be asking for 2 different cases?

jimthompson5910 (jim_thompson5910):

p = 0.25 (probability of success) q = 1-p (probability of failure) n = 20 (number of questions)

jimthompson5910 (jim_thompson5910):

let me think for a sec

OpenStudy (anonymous):

ok. so what about the introduction of the 5 students? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

"In a multiple choice there are 20 questions" there are 5 students. So there are really 5*20 = 100 questions. Agreed?

OpenStudy (anonymous):

huh.. alright...

OpenStudy (anonymous):

@jim_thompson5910 please continue :)

jimthompson5910 (jim_thompson5910):

sorry I'm thinking of a good way to word it

OpenStudy (anonymous):

@jim_thompson5910 oh sorry please take your time :3

jimthompson5910 (jim_thompson5910):

imagine you have 5 rows of boxes each row represents a person in each row you have 20 boxes (so you have 5*20 = 100 boxes total) each box represents a question. You can place C for correct in it, N for not correct.  I couldnt' fit all the 20 boxes in the drawing, so I used the dots (...) to represent it